a bullet loses 25% of its speed while travelling 14 cm inside a wooden block.how much further it travel before stopping?

Grade:12th pass

1 Answers

Shreyank
5 years ago
Let v=(100-25)m/s = 75m/s u=100m/s v¹=0m/s s = 14cm = 0.14m Then, a = (v² - u²)/2s =(75² - 100²)/2(0.14) a =-15625m/s² s¹ = ([v¹]² - v²)/2a s¹ = (0 - 75²)/2(-15625) s¹ = 0.18mTherefore the bullet will travel 0.18m further until it stops.